OpenStudy (anonymous):

Zarkon is it just 14 or -14?

OpenStudy (zarkon):

\[\sqrt{-196}=14i\] is that what you are talking about

OpenStudy (anonymous):

\[\sqrt{-196} = \sqrt{-1}\sqrt{196} = i\sqrt{196}\]

OpenStudy (anonymous):

yess but y did u put i?

OpenStudy (anonymous):

imaginary

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Complex_number

OpenStudy (anonymous):

Alchemista

OpenStudy (anonymous):

i need ur help

OpenStudy (zarkon):

if you don't just want the principle value then it is \[\pm\] the given answer.

OpenStudy (anonymous):

there is no real solution to \[\sqrt{-196}\] as -192<0 and -14^=196 and14^2=196 so the imaginary number i=sqroot-1 is used instead

OpenStudy (anonymous):

There is no real number whose square is a negative number. That's why the only solutions like on the complex field.

OpenStudy (anonymous):

so theres no solution to 196 sqaure root?

OpenStudy (anonymous):

Yes, a real solution to \(\sqrt{196}\) but not \(\sqrt{-196}\)

OpenStudy (zarkon):

the square root of 196 is \[\pm14\]

OpenStudy (anonymous):

yes i meant -196

OpenStudy (anonymous):

There are no real numbers whose square is a negative number. So for any negative number you will not have a real square root.

OpenStudy (anonymous):

remember thax x^2=x*x and never x*-x which would give a negative solution.

OpenStudy (anonymous):

Read about complex numbers.

OpenStudy (anonymous):

ok tanks